ln (xsquared + 3x + 4) = ln 2
Answer by stanbon(48569)

ln (xsquared + 3x + 4) = ln 2
Take the anti-log of both sides to get:
x^2+3x+4 = 2
x^2+3x+2=0
(x+2)(x+1)=0
x=-2 or x=-1
Checking in the original problem:
If x=-2: ln(4-6+4)=ln2
If x=-1: ln(1-3+4)=ln2
Both answers are good.
==============
Cheers,
Stan H.